# Detection Is Harder Than Estimation in Certain Regimes: Inference for Moment and Cumulant Tensors

### 27 Mar 2026 v1

@@ Abstract
We study estimation and detection of high-order moment and cumulant tensors from $n$ i.i.d. observations of a $p$-dimensional random vector, with performance measured in tensor spectral norm. On the statistical side, we prove that under sub-Gaussianity, the minimax rate for estimating the order-$d$ moment and cumulant tensors is $\sqrt{p/n}\wedge 1$. In contrast to covariance estimation, the sample moment tensor is generally no longer rate-optimal for higher-order moments. We therefore develop an estimator that attains the minimax rate up to logarithmic factors through a convex feasibility formulation over an $\varepsilon$-net of the unit sphere. On the computational side, we study the problem of testing whether the $d$-th order cumulant tensor vanishes after whitening. Using the low-degree polynomial framework, we provide evidence that detection is computationally hard when $n\ll p^{d/2}$. At the same time, we identify a regime in which an efficiently computable estimator attains error smaller than the separation at which low-degree tests can reliably distinguish the null from the alternative. This yields the striking conclusion that computationally efficient detection can be harder than computationally efficient estimation, revealing an unusual reverse detection–estimation gap: in a broad regime, computationally efficient estimation is possible at a smaller scale than computationally efficient detection. This phenomenon arises because the computational difficulty is driven not only by the statistical model, but also by the loss function itself: tensor spectral norm is NP-hard to compute. This feature makes the proposed open problems regarding computational lower bounds for estimation qualitatively different from the existing literature. Our results therefore uncover a new kind of computational–statistical gap.
@@ 1 Introduction
These two facts suggest a basic question: what is the computationally optimal way to estimate high-order moments and cumulants in tensor spectral norm? To understand computational barriers for estimation, we study an associated detection problem. Our results show that this question exhibits a rich computational–statistical structure that is qualitatively different from both covariance estimation and standard planted-signal problems. In particular, we identify a regime in which computationally efficient detection is harder than computationally efficient estimation.
@@
which is sharp in its dependence on $(n,p)$ even in the sub-Gaussian case $\beta=2$; see Section 3.2. Moreover, we construct a statistically optimal (up to a logarithmic factor) estimator via a convex feasibility formulation over an $\varepsilon$-net of the sphere, showing that the minimax benchmark $\sqrt{p/n}$ is statistically attainable; see Section 3.3. Finally, we show that detection and consistent estimation have the same statistical limit. This establishes the statistical landscape of the problem.
@@
$\sqrt{\frac{p^{d/2}}{n(\log(n))^{8}}}.$
@@
Comparing these two scales reveals an unusual phenomenon: the error of an efficiently computable estimator is asymptotically smaller than the separation at which low-degree tests can reliably distinguish the null from the alternative. In this sense, computationally efficient detection can be strictly harder than computationally efficient estimation. This is in contrast to the more familiar detection–recovery gap in planted problems, where recovery is typically the harder task (Bresler and Jiang, 2023; Mao et al., 2023; Schramm and Wein, 2022).
@@ 1.1 Related Work
In the ICA model, one observes a random vector of the form $X=AS$, where $A$ is an unknown deterministic mixing matrix and $S$ has entrywise independent coordinates. A key step in many ICA algorithms is to exploit fourth-order information, since after whitening, non-Gaussianity is often identified through deviations of the 4th moment (or 4th cumulant) from that of a Gaussian distribution. Consequently, both detection and estimation in ICA are intimately connected to the estimation of 4th moments and cumulants.
@@ 2 Preliminaries and Notation
We use $c,C,C^{\prime}$, etc. to denote positive constants whose values may vary from line to line, and use subscripts to indicate their dependence on parameters. For example, $C_{\beta,d}$ denotes a constant that depends only on $\beta$ and $d$, and not on any other parameters such as $p$. For a positive integer $d$, write $d]=\{1,\ldots,d\}$. Vectors are denoted by lowercase letters, matrices by uppercase letters, and tensors by calligraphic letters such as ${\mathbfcal{T}}\in\mathbb{R}^{p_{1}\times\cdots\times p_{d}}$. For $x\in\mathbb{R}^{p}$, let $\|x\|$ be its Euclidean norm, and let $\mathbb{S}^{p-1}=\{x\in\mathbb{R}^{p}:\|x\|=1\}$ be the unit sphere. For any random variable or random vector $X$, $Y$, we write $X\overset{d}{\sim}Y$ if they have the same distribution. For a matrix $A$, let $\|A\|$ and $\|A\|_{F}$ denote its spectral norm and Frobenius norm, respectively. For a tensor ${\mathbfcal{T}}\in(\mathbb{R}^{p})^{\otimes d}$, its spectral norm is
@@ 4.2 Detection Lower Bound
be the whitening matrix for ${\mathbfcal{K}}_{2}(X)$. Conditioning on $U$, the first d1d-1 cumulants of the transformed random vector $SX$ coincide with those of a standard Gaussian distribution, while
@@
Thus, for any fixed constant $a>0$ and $\|U\|=\sqrt{p}$, the spectral norm of the order-$d$ cumulant of $SX$ is bounded below by a positive constant $C_{d}>0$ depending only on $d$.
@@ 4.3 A “Reverse” Detection–Estimation Gap
$\|{\mathbfcal{K}}_{d}(SX)\|=c_{d}\left(\frac{a}{1+a}\right)^{d/2}\gtrsim\sqrt{\frac{p^{d/2}}{n(\log(n))^{8}}}:=D_{\operatorname{det}}.$
@@ B.2 Proof of Theorem 3
Because $h_{M}$ is bounded, $A$ is analytic near 0, so
@@ B.5 Proof of Theorem 6
The first statement, i.e., the high probability bound for $\widehat{\mathbfcal{M}}$, immediately follows from the discussion in Section 3.3. The statement about expectation follows as $\|\widehat{\mathbfcal{M}}\|$ and $\|\mathbb{E}X^{\otimes d}\|$ are both upper bounded. Finally, for cumulant tensor estimation, by (1) and the fact that $F_{d}$ is a fixed tensor polynomial in the lower-order moments, with degree depending only on $d$, the results follow.
@@ B.6 Proof of Lemma 1
We next calculate $\|L_{\leq D}\|^{2}:=\|L_{\leq D}\|^{2}_{L^{2}(H_{0})}$. We use the following notations: for vector $\alpha=(\alpha_{1},\ldots,\alpha_{m}),\alpha_{k}\in\mathbb{N}$, denote $|\alpha|=\sum_{k\in[m]}\alpha_{k}$, $\alpha!=\prod_{k\in[m]}\alpha_{k}!$, and multivariate Hermite polynomials $H_{\alpha}(x)=\prod_{k\in[m]}H_{\alpha_{i}}(x_{i})$. For vector $u\in\mathbb{R}^{m}$, let $u^{\alpha}=\prod_{k\in[m]}u_{k}^{\alpha_{k}}$. For a $np$-dimensional vector $y$, we use denote $y=(y^{(1)},\ldots,y^{(n)})^{\top}$ with $y^{(m)}\in\mathbb{R}^{p}$.

### 04 Apr 2026 v2

@@ Abstract
We study estimation and detection of high-order moment and cumulant tensors from $n$ i.i.d. observations of a $p$-dimensional random vector, with performance measured in tensor spectral norm. Under sub-Gaussianity, we show that the minimax rate for estimating the order-$d$ moment and cumulant tensors is $\sqrt{p/n}\wedge 1$. In contrast to covariance estimation, the sample moment tensor is generally not rate-optimal for $d\geq 3$, and we construct an estimator that attains the minimax rate up to logarithmic factors. On the computational side, we study testing whether the $d$-th order cumulant tensor vanishes after whitening. Using the low-degree polynomial framework, we provide evidence that detection is computationally hard when $n\ll p^{d/2}$. At the same time, we identify a regime in which an efficiently computable estimator achieves error smaller than the separation at which low-degree tests can reliably distinguish the null from the alternative. This reveals an unusual reverse detection–estimation gap: computationally efficient detection can be harder than computationally efficient estimation. The underlying reason is that the relevant loss, tensor spectral norm, is itself NP-hard to compute, creating a new form of computational–statistical gap.
@@ 1 Introduction
These two facts motivate a basic question: what is the computationally optimal approach to estimating high-order moments and cumulants in tensor spectral norm? To investigate the computational barriers to estimation, we study an associated detection problem. Our results reveal a rich computational–statistical structure that differs qualitatively from both covariance estimation and more standard planted-signal problems. In particular, this phenomenon is distinct from the more familiar detection–recovery gap in the literature, where recovery is typically the harder task Schramm and Wein (2022); Luo and Gao (2024); Mao et al. (2023); Bresler and Jiang (2023); see Sections 1.1 and 4.3 for further discussion.
@@
which is sharp in its dependence on $(n,p)$ even in the sub-Gaussian case $\beta=2$; see Section 3.2. Moreover, we construct a statistically optimal (up to a logarithmic factor) estimator via a convex feasibility formulation over an $\varepsilon$-net of the sphere, showing that the minimax benchmark $\sqrt{p/n}$ is statistically attainable; see Section $3.$ Finally, we show that detection and consistent estimation have the same statistical limit. This establishes the statistical landscape of the problem.
@@
$\sqrt{\frac{p^{d/2}}{n\operatorname{polylog}(n)}}.$
@@
Comparing these two scales reveals an unusual phenomenon: the error of an efficiently computable estimator is asymptotically smaller than the separation at which low-degree tests can reliably distinguish the null from the alternative. In this sense, computationally efficient detection can be strictly harder than computationally efficient estimation. This is in contrast to the more familiar detection–recovery gap in planted problems, where recovery is typically the harder task.
@@ 1.1 Related Work
Detection–estimation gap
The relationship among detection, estimation, and recovery has long been studied from both classical statistical and more recent computational perspectives. For example, Ingster et al. (2010) develops sharp minimax detection boundaries in sparse regression and shows that procedures based on estimating the regression coefficients in $\ell_{2}$ norm generally require stronger signals than those needed for optimal detection. In sparse heterogeneous mixtures, Donoho and Jin (2004) characterizes the sharp detection boundary in the rare/weak regime and shows that higher criticism attains this boundary adaptively, thereby illustrating that global detection can succeed even when the underlying sparse signals are too weak for reliable identification of the affected coordinates. In network models, Decelle et al. (2011) predicts detectability phase transitions in the stochastic block model, while Abbe et al. (2016); Abbe and Sandon (2015) establishes sharp thresholds for exact and partial recovery, again highlighting that different inferential goals can have different thresholds. More recently, Bresler and Jiang (2023); Mao et al. (2023) studies computational gaps between detection and stronger inferential tasks, including recovery and refutation, in planted-signal problems. Most of the works concern settings in which recovery or estimation is harder than detection; by contrast, our setting exhibits the opposite phenomenon, where detection can be strictly harder than estimation.
@@
In the ICA model, one observes a random vector of the form $X=AS$, where AA is an unknown deterministic mixing matrix and $S$ has entrywise independent coordinates. A key step in many ICA algorithms is to exploit fourth-order information, since after whitening, non-Gaussianity is often identified through deviations of the 4th moment (or 4th cumulant) from that of a Gaussian distribution. Consequently, both detection and estimation in ICA are intimately connected to the estimation of 4th moments and cumulants.
@@ 2 Preliminaries and Notation
We use $c,C,C^{\prime}$, etc. to denote positive constants whose values may vary from line to line, and use subscripts to indicate their dependence on parameters. For example, $C_{\beta,d}$ denotes a constant that depends only on $\beta$ and $d$, and not on any other parameters such as pp. For a positive integer $d$, write $d]=\{1,\ldots,d\}$. Vectors are denoted by lowercase letters, matrices by uppercase letters, and tensors by calligraphic letters such as ${\mathbfcal{T}}\in\mathbb{R}^{p_{1}\times\cdots\times p_{d}}$. For $x\in\mathbb{R}^{p}$, let $\|x\|$ be its Euclidean norm, and let $\mathbb{S}^{p-1}=\{x\in\mathbb{R}^{p}:\|x\|=1\}$ be the unit sphere. For any random variable or random vector $X$, $Y$, we write $X\overset{d}{\sim}Y$ if they have the same distribution. For a matrix AA, let $\|A\|$ and $\|A\|_{F}$ denote its spectral norm and Frobenius norm, respectively. For a tensor ${\mathbfcal{T}}\in(\mathbb{R}^{p})^{\otimes d}$, its spectral norm is
@@ 4.2 Detection Lower Bound
be the whitening matrix for ${\mathbfcal{K}}_{2}(X)$. Conditioning on $U$, the first d1d-1 cumulants of the transformed random vector SXSX coincide with those of a standard Gaussian distribution, while
@@
Thus, for any fixed constant $a>0$ and $\|U\|=\sqrt{p}$, the spectral norm of the order-$d$ cumulant of SXSX is bounded below by a positive constant $C_{d}>0$ depending only on $d$.
@@ 4.3 A “Reverse” Detection–Estimation Gap
$\|{\mathbfcal{K}}_{d}(SX)\|=c_{d}\left(\frac{a}{1+a}\right)^{d/2}\gtrsim\sqrt{\frac{p^{d/2}}{n(\log(n))^{8d}}}:=D_{\operatorname{det}}.$
@@ B.2 Proof of Theorem 3
Because $h_{M}$ is bounded, AA is analytic near 0, so
@@ B.5 Proof of Theorem 6
The first statement, i.e., the high probability bound for $\widehat{\mathbfcal{M}}$, immediately follows from the discussion in Section $3.$ The statement about expectation follows as $\|\widehat{\mathbfcal{M}}\|$ and $\|\mathbb{E}X^{\otimes d}\|$ are both upper bounded. Finally, for cumulant tensor estimation, by (1) and the fact that $F_{d}$ is a fixed tensor polynomial in the lower-order moments, with degree depending only on $d$, the results follow.
@@ B.6 Proof of Lemma 1
We next calculate $\|L_{\leq D}\|^{2}:=\|L_{\leq D}\|^{2}_{L^{2}(H_{0})}$. We use the following notations: for vector $\alpha=(\alpha_{1},\ldots,\alpha_{m}),\alpha_{k}\in\mathbb{N}$, denote $|\alpha|=\sum_{k\in[m]}\alpha_{k}$, $\alpha!=\prod_{k\in[m]}\alpha_{k}!$, and multivariate Hermite polynomials $H_{\alpha}(x)=\prod_{k\in[m]}H_{\alpha_{i}}(x_{i})$. For vector $u\in\mathbb{R}^{m}$, let $u^{\alpha}=\prod_{k\in[m]}u_{k}^{\alpha_{k}}$. For a npnp-dimensional vector $y$, we use denote $y=(y^{(1)},\ldots,y^{(n)})^{\top}$ with $y^{(m)}\in\mathbb{R}^{p}$.